OpenStudy (cruffo):

solve for x x-3*cuberoot(x) = 0

OpenStudy (cruffo):

I was using wolfram alpha to graph, and it's only showing two of three roots. It seems to think the domain is only x larger or equal to 0.

OpenStudy (anonymous):

Is it (x-3)*cuberoot(x) = 0?

OpenStudy (cruffo):

sorry, \[x - 3\sqrt[3]{x}\]

OpenStudy (cruffo):

the graph on wolfram does not match the information I get from looking at intervals of increasing, decreasing, local extrema, concavity, domain, ...

OpenStudy (anonymous):

It's going to be difficult to isolate x here

OpenStudy (cruffo):

factor out the cube root \[\sqrt[3]{x}(\sqrt[3]{x^2} - 3)\]

OpenStudy (anonymous):

You gave me the impression that the x -3 wasn't in parentheses

myininaya (myininaya):

so you are trying to figure out why wolfram has a different graph then the one you came up with?

OpenStudy (cruffo):

it isn't...

OpenStudy (anonymous):

Even so, you still have an x lingering

OpenStudy (cruffo):

myininaya: yes - exactly

myininaya (myininaya):

ok let me see what i come up with i will scan and post ok?

OpenStudy (cruffo):

thanks... I attached what I

OpenStudy (cruffo):

have

OpenStudy (anonymous):

there are three x values that work, but only two solutions are valid

OpenStudy (cruffo):

why is the negative sol. not valid?

OpenStudy (anonymous):

x = 0, 3sqrt 3, -3sqrt 3

OpenStudy (cruffo):

heromiles, \[0-3\sqrt[3]{0} = 0\]

OpenStudy (cruffo):

\[-3\sqrt[3]{3\sqrt{3}}= -3\sqrt[6]{27} = -3(3^3)^{1/6} = -3\sqrt{3} \] and \[-3\sqrt[3]{-3\sqrt{3}} = 3\sqrt[6]{27} = 3(3^3)^{1/6} = 3\sqrt{3} \]

OpenStudy (anonymous):

What about the graph? Does it match your solutions?

myininaya (myininaya):

your graph looks fine

OpenStudy (cruffo):

the TI-84 Plus matches, but not wolfram alpha

OpenStudy (cruffo):

Thanks for checking myininaya. Any thoughts on why wolfram disagrees?

myininaya (myininaya):

no i dont know its weird i was like maybe i need to zoom in but no wolfram has the wrong graph

myininaya (myininaya):

even my calculator agrees with us

OpenStudy (anonymous):

You graphed two solutions, right?

myininaya (myininaya):

you mean two x-intercepts? that doesn't matter i still have the same shape

OpenStudy (anonymous):

Yes, two x-intercepts

OpenStudy (anonymous):

I solved it a bit differently though

myininaya (myininaya):

did you factor out an x instead?

OpenStudy (cruffo):

myininaya: I think your second derivative should be \[\frac{2}{3x^{5/3}} = \frac{2}{3x\sqrt[3]{x^2}}\] So there would be a sign change at 0, right?

OpenStudy (anonymous):

No, I cubed both sides to get rid of the cube root

myininaya (myininaya):

you are right cruffo

OpenStudy (anonymous):

I ended up with x(x^2 - 27) = 0

myininaya (myininaya):

but still this doesn't give us the graph wolfram has

myininaya (myininaya):

what is the point of the second derivative lol

OpenStudy (cruffo):

concavity :)

myininaya (myininaya):

i know it tells if the function is concave up and concave down lol but it doesn't really change the graph

myininaya (myininaya):

very pretty

OpenStudy (cruffo):

no :), of course. 2nd derivative just gives us a little more information to work with.

OpenStudy (anonymous):

I thought the original question was to find x.... I simply used algebra to find x, then I plugged in the values of x and found that x = 0, sqrt 27 Then I graphed it. What more do you need to do?

myininaya (myininaya):

he wants to know why his graph is different from wolfram

OpenStudy (cruffo):

so do I :)

myininaya (myininaya):

you are totally right and wolfram is wrong

OpenStudy (cruffo):

I wanted to assign this problem for homework, but if wolfram is gonna be that way about it...

OpenStudy (anonymous):

Cruffo, you graphed it manually?

OpenStudy (cruffo):

yep

OpenStudy (cruffo):

Thank you all for your help.

myininaya (myininaya):

wait cruffo

myininaya (myininaya):

wolfram is confused about the cube root of -1 i think it thinks it imagary do -1-cuberoot(-1)

OpenStudy (cruffo):

wow!!!

myininaya (myininaya):

or i mean do -1-3cuberoot(-1)

OpenStudy (cruffo):

cube roots of unity?????

myininaya (myininaya):

both should be imagary based off wolfram let me try

myininaya (myininaya):

yep

OpenStudy (cruffo):

anyway to force it to work only over the reals...

myininaya (myininaya):

cuberoot(-1)=0.5+ci where c is some irrational number

OpenStudy (anonymous):

Why are you so dependent on wolfram alpha for graphing?

OpenStudy (cruffo):

I'm not, but my students like to use it

myininaya (myininaya):

he proabably don't want his students second guessing themselves since there are not confident in math

OpenStudy (cruffo):

Very correct (and it's she by the way :), and I try to help them with the correct commands to use. I don't mind that they use a computer, I like it. In fact, I prefer them to use a computer to a calculator. But, this will cause the majority of them to give up. I can't seem to figure out how to force wolfram to work only on the real number without errors...

myininaya (myininaya):

me neither

OpenStudy (cruffo):

oh well. I'll just give the warning.

myininaya (myininaya):

i'm a she too just so everyone knows

OpenStudy (anonymous):

Contact Wolfram and point out the error

myininaya (myininaya):

so don't call me sir or boy

myininaya (myininaya):

lol

OpenStudy (cruffo):

Yes!!! :) Again, thank you all for your help.

myininaya (myininaya):

this doesn't make sense x^3=-1 wolfram gives the solution x=-1

myininaya (myininaya):

i mean its right but it can't do (-1)^(1/3) so how can it solve this

myininaya (myininaya):

i wonder what method it used to solve x^3=-1

OpenStudy (anonymous):

Hilarious

OpenStudy (anonymous):

For graphing , I prefer my TI N-spire

OpenStudy (cruffo):

I'll send an email regarding the issue to wolfram, along with these other test. I'm not too familiar with Mathematica, but I think it works "symbolically", but not sure... I prefer SAGE. Very nice open source program for symbolic math. You can specify the field you want to work on. http://www.sagemath.org/

myininaya (myininaya):

ok i don't trust wolfram anymore

myininaya (myininaya):

thanks for letting me know about this lol

OpenStudy (anonymous):

Geogebra is a pretty good program

OpenStudy (cruffo):

Well, you should always try to verify answers at least qualitatively, right. This is actually a really good example of that fact! I always try to find those sneaky functions to use, but this one crept up on me :) Ohh... I like geogebra!!! I should have though about that one.

myininaya (myininaya):

we built programs its our fault they make mistakes

OpenStudy (cruffo):

Geogebra does a good graph of this problem

OpenStudy (anonymous):

I graphed it on geogebra...it graphed correctly and it finds the cube root of -1 correctly

myininaya (myininaya):

nice

myininaya (myininaya):

lets find a mistake in the program lol

myininaya (myininaya):

i will save it for later i need to go to bed

OpenStudy (anonymous):

The only question is, can it isolate and solve for x

OpenStudy (cruffo):

good night myininaya

OpenStudy (cruffo):

not sure, heromiles. I haven't used geogebra for that.

OpenStudy (anonymous):

If there's a way, I'll figure it out

OpenStudy (cruffo):

there is a "root" command, but not a "solve" command

OpenStudy (cruffo):

root command uses newton's and false position methods

OpenStudy (cruffo):

it finds the positive and negative root, but not the root at 0 using newton's method. It will find the root at 0 using false position.

OpenStudy (cruffo):

well, I need to finish up the hints for the homework problems. Good night!!

OpenStudy (anonymous):

There's also a solutions command apparently. Have to try it out.

OpenStudy (anonymous):

Well, I think this is only for version 4.0

OpenStudy (anonymous):

Actually, it has a solutions and solve command, which is great.